OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2)

OpenStudy (anonymous):

first find the slope from these two points

OpenStudy (diyaja):

put it in the eqn (y-y1)=m(x-x1), where m is the gradient

OpenStudy (diyaja):

m=(y2-y1)/(x2-x1)